Please compute cost per unit of each product on the new method of allocation The information related to manufacturing is given below. WRITERS' Company Year 2018-19 Fountain Pen Ball point Units 10000 20000 Area occupied (sq feet) 7000 8000 Variable cost – Direct 32 25 Fixed Cost - Direct (per unit) 10 8 Rent - Fixed Cost per month (Indirect - common for all products) 3,00,000.00
1
Expert's answer
2020-12-23T07:38:46-0500

Total variable cost = "variable cost per unit \times units" = "Rs 32\times 10000 units" = "Rs 320,000"

Total fixed cost = "fixed cost per unit \times units = Rs 10 \times 10000 units = Rs 100,000"

Total Fountain’s rent = 

Total products’ units = 10000 + 20000 = 30000

Total Fountain’s rent = "\frac{10000}{30000} \times Rs 300,000 = Rs 100,000"

Total cost = total variable cost + total fixed cost + total rent

= Rs 320,000 + Rs 100,000 + Rs 100,000

= Rs 520,000

Cost per unit = "\frac{Total cost}{total units} = \frac{ Rs 520,000}{10000 units } =Rs 52"


Pen ball point’s cost per unit

Total variable cost = "variable cost per unit \\times \tunits = Rs 25 \\times 20000 units = Rs 500,000"

"Total fixed cost = fixed cost per unit \times units = Rs 8 \times 20000 units = Rs 160,000"

Total Pen ball point’s rent = 

"\frac{ Pen ball point’s units }{ total products’ units } \times the total fixed indirect cost of rent"

Total products’ units = 10000 + 20000 = 30000

Total Pen ball point’s rent = "\frac{ 20000}{ 30000} \times Rs 300,000 = Rs 200,000"

Total cost = total variable cost + total fixed cost + total rent

= Rs 500,000 + Rs 160,000 + Rs 200,000

= Rs 860,000

"Cost per unit = \frac{Total cost }{ total units } = \frac{ Rs 860,000}{ 20000 units } = Rs 43"